# Variational inequalities and smooth-fit principle for singular stochastic control problems in Hilbert spaces

### v1

@@ 4. Selecting a specific direction of action: A related optimal stopping problem
Let $(I^{k})_{k\in\mathbb{N}}:=(\hat{n},\nu^{k})_{k\in\mathbb{N}}\subseteq\mathcal{I}_{0}$ be a minimizing sequence for $V(x)$. Let us denote by $X^{x,k}:=X^{x,I^{k}}$, $k\in\mathbb{N}$. Given that $V(x)\leq\hat{c}_{o}(1+|x|^{p}_{H})$, $x\in H$, (cf. Proposition 3.1) for $p\geq 2$ as in Assumption 2.8, we have for some $\varepsilon>0$
(4.4) $\sup_{k\in\mathbb{N}}\mathsf{E}\bigg{[}\int_{0^{-}}^{\infty}e^{-\rho t}|X^{x,k}_{t}|_{H}^{p}\,\mathrm{d}t\bigg{]}\leq\frac{1}{\kappa_{1}}\Big{(}\hat{c}_{o}(1+|x|^{p}_{H})+\varepsilon+\frac{\kappa_{2}}{\rho}\Big{)}.$
@@
(4.5) $|\widehat{n}|_{H}^{p}e^{\lambda tp}\mathsf{E}\big{[}|\nu^{k}_{t}|^{p}\big{]}\leq C\Big{(}|x|_{H}^{p}+\overline{c}_{p}+\mathsf{E}\big{[}|X^{x,k}_{t}|^{p}\big{]}\Big{)}.$
@@
(4.6) $\sup_{k\in\mathbb{N}}\mathsf{E}\bigg{[}\int_{0}^{\infty}e^{-(\rho-\lambda p)t}|\nu^{k}_{t}|^{p}\mathrm{d}t\bigg{]}\leq C\Big{(}1+|x|_{H}^{p}+\sup_{k\in\mathbb{N}}\mathsf{E}\bigg{[}\int_{0}^{\infty}e^{-\rho t}|X^{x,k}_{t}|^{p}\mathrm{d}t\bigg{]}\Big{)}<\infty.$
Hence, by Banach-Saks theorem, there exist a subsequence of $(\nu^{k})_{k\in\mathbb{N}}$ still denoted by $(\nu^{k})_{k\in\mathbb{N}}$ and some ${\nu}^{\star}\in L^{p}(\Omega\times[0,\infty);\mathsf{P}\otimes e^{-(\rho-\lambda p)t}\mathrm{d}t)$ such that
(4.7) $\widehat{\nu}^{j}:=\frac{1}{j+1}\sum_{k=0}^{j}\nu^{k}\to{\nu}^{\star}\quad\text{in}\,\,L^{p}(\Omega\times[0,\infty);\mathsf{P}\otimes e^{-(\rho-\lambda p)t}\mathrm{d}t).$
Actually, up to passing to a further subsequence (again, relabeled in the following), such a convergence can be realized $\mathsf{P}\otimes e^{-(\rho-\lambda p)t}\mathrm{d}t$-a.e. Then, by arguing as in [44, Lemmata 4.5-4.7], the process $\nu^{\star}$ admits a modification - still denoted by $\nu^{\star}$ that is right-continuous, nondecreasing and $\mathbb{F}$-adapted, and thus belongs to $\mathcal{S}$. Furthermore, given that ${\nu}^{\star}\in L^{p}(\Omega\times[0,\infty);\mathsf{P}\otimes e^{-(\rho-\lambda p)t}\mathrm{d}t)$ and that $t\mapsto\mathsf{E}[|\nu^{\star}_{t}|^{p}]$ is nondecreasing, it follows that the integrability condition required in (2.17) is also met (recall that, in this section, $\vartheta_{t}\equiv\widehat{n}$). Set then $\widehat{I}^{j}:=(\hat{n},\widehat{\nu}^{j})\in\mathcal{I}_{0}$, $j\in\mathbb{N}$, and $I^{\star}:=(\hat{n},\nu^{\star})\in\mathcal{I}_{0}$.
@@
The process $\xi\in\mathcal{I}_{0}$ and it has direction of action $\hat{n}$ and intensity of action $\nu^{\xi}$ such that $\nu^{\xi}_{0^{-}}=0$ and $\mathrm{d}\nu^{\xi}_{t}=\mathrm{d}\nu^{\star}_{t}+\varepsilon e^{\lambda t}\delta(t-\tau)$, for any $t\geq 0$. Furthermore, letting $X^{x-\varepsilon\hat{n},\xi}$ be the solution to (2.7) started at time $0^{-}$ from level $x-\varepsilon\hat{n}$ and controlled by $\xi$, we have that: $X^{x-\varepsilon\hat{n},\xi}\equiv X^{x-\varepsilon\hat{n},I^{\star}}$ on $[0,\tau)$, while $X^{x-\varepsilon\hat{n},\xi}\equiv X^{x,I^{\star}}$ on $[\tau,\infty)$.
@@
and notice that denoting by $X^{x+\varepsilon\hat{n},\eta}$ the solution to (2.7) started at time $0^{-}$ from level $x+\varepsilon\hat{n}$ and controlled by $\eta$, we have that: $X^{x+\varepsilon\hat{n},\eta}\equiv X^{x+\varepsilon\hat{n},0}$ on $[0,\tau^{\varepsilon})$, while $X^{x+\varepsilon\hat{n},\eta}\equiv X^{x,I^{\star}}$ on $[\tau^{\varepsilon},\infty)$.
@@
By continuity of $U$, the stopping region $\{x\in H:\,U(x)=\Phi(x)\}$ is closed. Hence, by standard theory of optimal stopping (see, e.g., [55, Ch. I, Sec. 2.2], whose results hold for an underlying process taking values in a measurable space), one has $\mathsf{P}$-a.s. that
@@
(4.18) $U(x)=\sup_{\tau\in\mathcal{T}}\mathsf{E}\left[e^{-(\rho-\lambda)\tau}\mathds{1}_{{\tau<\theta}}\Phi(X^{x}_{\tau})+e^{-(\rho-\lambda)\theta}\mathds{1}_{{\tau\geq\theta}}U(X^{x}_{\theta})\right],\ \ \ \ \forall x\in H.$
We refrain from providing a proof of the previous result. As a matter of fact, this would follow the same lines and steps as in the finite-dimensional settings (see, e.g., [56, Sec. 5.2] and references therein) and exploit the flow property of $X$, which is ensured here by the semigroup property.
@@
with $\mathcal{G}$ as in (3.3), and for which, recalling the class (3.2), we now provide the definition of viscosity solution.
@@
    We say that $u\in C(H)$ is a viscosity supersolution to (4.19) at $x\in H$ if, for every $\varphi\in\mathcal{X}$ such that $0=u(x)-\varphi(x)=\min(u-\varphi)$, one has
    $\min\big{\{}((\rho-\lambda)-\mathcal{G})\varphi,\ \varphi-\Phi\big{\}}\geq 0.$
@@
    We say that $u\in C(H)$ is a viscosity subsolution to (4.19) at $x\in H$ if, for every $\varphi\in\mathcal{X}$ such that $0=u(x)-\varphi(x)=\max(u-\varphi)$, one has
    $\min\big{\{}((\rho-\lambda)-\mathcal{G})\varphi,\ \varphi-\Phi\big{\}}\leq 0.$
@@
with $\mathcal{B}_{|\cdot|_{H}}(x,\varepsilon):=\{y\in H:\,|y-x|_{H}\leq\varepsilon\}.$ There exists $m_{o}>0$ such that
@@
We follow the ideas of the proof in the finite-dimensional setting proposed by [56, Thm. 5.2.1], but we simplify substantially the proof of the subsolution property thanks to Lemma 4.8.
Supersolution property. Recall (3.2) and let $\varphi\in\mathcal{X}$ be such that
@@
By continuity,
$((\rho-\lambda)-\mathcal{G})\varphi(y)>\eta\ \ \mbox{and}\ \ U(y)-\Phi(y)>\frac{\eta}{\rho-\lambda}\ \ \ \forall\mathcal{B}_{|\cdot|_{H}}(x,\varepsilon),$
where we recall that $\mathcal{B}_{|\cdot|_{H}}(x,\varepsilon):=\{y\in H:\,|y-x|_{H}\leq\varepsilon\}.$
@@
Then, using Dynkin’s formula of Proposition 3.3 (applied again in the case of $I$ being the null control) we obtain for each $\tau\in\mathcal{T}$
@@
If one aims at a comparison theorem for (4.19), this might be proved by adapting the techniques of [36] in order to deal with the unbounded operator in the dynamics of the state process (not present in [36]). To that end, one should treat the unbounded term as in the regular control case by adding radial functions as test functions in the definition of viscosity solutions (see [27, Ch. 3]).
@@
so that ${G}_{\hat{n}}$ is clearly concave, belongs to $C^{1}(H)$ and it has sublinear growth with respect to $|\cdot|_{H}$. Moreover, in the first case, setting $k:=\mathcal{A}h$, one has
@@
Using (4.1) and that $|e^{t\mathcal{A}}|_{\mathcal{L}(H)}\leq e^{-\delta t}$, one easily finds from (4.15) that the semiconcavity with respect to the $|\cdot|_{-1}$ norm of ${G}_{\hat{n}}$ implies semiconvexity of $\Phi$ with respect to the $|\cdot|_{-1}$ norm as well. Hence, given that $x\mapsto X^{x,0}_{\cdot}$ is linear, we find that $U$ is semiconvex with respect to the $|\cdot|_{-1}$ norm being supremum of semiconvex functions (see [14, Prop. 2.1.5], whose proof does not suffer the dimensionality of the state space).
By (4.15), one also has that the growth condition on ${G}_{\hat{n}}$, together with (4.23), imply that $|\Phi(x)|\leq K_{\Phi}(1+|x|_{-1})$, for a suitable $K_{\Phi}>0$. The latter, together with (4.24), in turn gives for $\widehat{\kappa}_{o}>0$
@@
Finally, the last claim is due to [25, Cor. 2.4, p. 12] and to the fact that, being $U$ semiconvex with respect to $|\cdot|_{-1}$ norm, by Lemma (i)(i) one can write
@@
In the rest of the paper, we are going to study the regularity properties of the (sub)gradient of $U$.
@@
The following hold true:
  1. (i)
    $D^{-}U(x)\subseteq\mathcal{D}(\mathcal{A})$ at all $x\in H$;
  2. (ii)
[…]
@@
Recall that by (i) in Lemma (i) we can write
@@
$D^{-}U(x)=D^{-}U_{0}(x)-\widehat{\kappa}_{1}\mathcal{A}^{-1}x.$
Hence, without loss of generality, up to replace $U$ by $U_{0}$, we work in the rest of this proof under the assumption that $U$ is convex. We now prove each item separately.
Proof of (i). Let $x\in H$, $(h_{n})\subseteq\mathcal{D}(\mathcal{A})$ be such that $h_{n}\to 0$ with respect to $|\cdot|_{H}$, and set $\hat{h}_{n}:=\mathcal{A}h_{n}$, so that also $|\hat{h}_{n}|_{-1}\to 0$. For $p\in D^{-}U(x)$, by convexity of $U$ we have
@@
is $|\cdot|_{H}$-upper semicontinuous by (4.25); that is, $\limsup_{n\rightarrow\infty}T_{x}(h_{n})\leq 0$. Taking now the sequence $(-h_{n})\subseteq\mathcal{D}(\mathcal{A})$ and exploiting the linearity of $T_{x}$, it also holds that $T_{x}$ is $|\cdot|_{H}$-lower semicontinuous. Hence, it is $|\cdot|_{H}$-continuous, which means that $p\in\mathcal{D}(\mathcal{A})$.
Proof of (ii). Consider the function
$u:\mathcal{D}(\mathcal{A})\to\mathbb{R},\ \ \ \ u(x):=U(\mathcal{A}x),\quad x\in H,$
so that $U=u\circ\mathcal{A}^{-1}.$ Then $u$ is convex and by Proposition 4.13 one has
$|u(x)|=|U(\mathcal{A}x)|\leq\widehat{\kappa}_{o}(1+|\mathcal{A}x|_{-1})=\widehat{\kappa}_{o}(1+|\mathcal{A}^{-1}\mathcal{A}x|_{H})=\widehat{\kappa}_{o}(1+|x|_{H}),\quad x\in H.$
Hence, $u$ is $|\cdot|_{H}$-locally Lipschitz continuous by [25, Cor. 2.4, p. 12] and thus it can be extended to a continuous function defined to the whole space $H$. Moreover, for all $x\in\mathcal{D}(\mathcal{A})$, we have
$q\in D^{-}u(x)\ \Longleftrightarrow\ p=\mathcal{A}q\in D^{-}U(\mathcal{A}x).$
Hence, by using the definition of subgradient,
$D^{-}u(x)=\mathcal{A}DU^{-}(\mathcal{A}x),\ \ \ \ \forall x\in\mathcal{D}(\mathcal{A}).$
By convexity, the multivalued map
$(\mathcal{D}(\mathcal{A}),|\cdot|_{H})\to(H,|\cdot|_{H}),\ \ x\mapsto D^{-}u(x)$
is upper hemicontinuous. Because
$D^{-}U(x)=\mathcal{A}^{-1}D^{-}u(\mathcal{A}^{-1}x),$

### v2

@@ 4. Selecting a specific direction of action: A related optimal stopping problem
Remark 4.3.
Within this section, the decision maker can act only on the intensity of the control, which is modeled as a one-dimensional singular control process belonging to the set $\mathcal{S}$. This assumption is made in order to develop regularity results for the value function $V$ by linking it to an optimal stopping problem. Indeed, the connection between singular stochastic control and optimal stopping is well understood for one-dimensional control processes of bounded variation, whereas it has not yet been established for multi-dimensional controls. Moreover, when the dimensionality of the problem is greater than one (but still finite) and the control is multi-dimensional, analyzing the geometry of the state space and constructing an optimal control rule of reflecting type already present significant challenges; see the introduction of [24] for a discussion of this point. These challenges become even more substantial in infinite-dimensional settings, where it is not even clear what precise meaning should be assigned to a “reflecting policy.” These highly intriguing aspects remain important open questions for future research.
@@
Let $(I^{k})_{k\in\mathbb{N}}:=(\hat{n},\nu^{k})_{k\in\mathbb{N}}\subseteq\mathcal{I}_{0}$ be a minimizing sequence for $V(x)$. Let us denote by $X^{x,k}:=X^{x,I^{k}}$, $k\in\mathbb{N}$. Given that $V(x)\leq\hat{c}_{o}(1+|x|^{2}_{H})$, $x\in H$, (cf. Proposition 3.1), we have for some $\varepsilon>0$
(4.4) $\sup_{k\in\mathbb{N}}\mathsf{E}\bigg[\int_{0^{-}}^{\infty}e^{-\rho t}|X^{x,k}_{t}|_{H}^{2}\,\mathrm{d}t\bigg]\leq\frac{1}{\kappa_{1}}\Big(\hat{c}_{o}(1+|x|^{2}_{H})+\varepsilon+\frac{\kappa_{2}}{\rho}\Big).$
@@
(4.5) $|\widehat{n}|_{H}^{2}e^{2\lambda t}\mathsf{E}\big[|\nu^{k}_{t}|^{2}\big]\leq C\Big(|x|_{H}^{2}+\overline{c}_{2}+\mathsf{E}\big[|X^{x,k}_{t}|^{2}\big]\Big).$
@@
(4.6) $\sup_{k\in\mathbb{N}}\mathsf{E}\bigg[\int_{0}^{\infty}e^{-(\rho-2\lambda)t}|\nu^{k}_{t}|^{2}\mathrm{d}t\bigg]\leq C\Big(1+|x|_{H}^{2}+\sup_{k\in\mathbb{N}}\mathsf{E}\bigg[\int_{0}^{\infty}e^{-\rho t}|X^{x,k}_{t}|^{2}\mathrm{d}t\bigg]\Big)<\infty.$
Hence, by Banach-Saks theorem, there exist a subsequence of $(\nu^{k})_{k\in\mathbb{N}}$ - still denoted by $(\nu^{k})_{k\in\mathbb{N}}$ - and some ${\nu}^{\star}\in L^{2}(\Omega\times[0,\infty);\mathsf{P}\otimes e^{-(\rho-2\lambda)t}\mathrm{d}t)$ such that
(4.7) $\widehat{\nu}^{j}:=\frac{1}{j+1}\sum_{k=0}^{j}\nu^{k}\to{\nu}^{\star}\quad\text{in}\,\,L^{2}(\Omega\times[0,\infty);\mathsf{P}\otimes e^{-(\rho-2\lambda)t}\mathrm{d}t).$
Actually, up to passing to a further subsequence (again, relabeled in the following), such a convergence can be realized $\mathsf{P}\otimes e^{-(\rho-2\lambda)t}\mathrm{d}t$-a.e. Then, by arguing as in [45, Lemmata 4.5-4.7], the process $\nu^{\star}$ admits a modification - still denoted by $\nu^{\star}$ - that is right-continuous, nondecreasing and $\mathbb{F}$-adapted, and thus belongs to $\mathcal{S}$. Furthermore, given that ${\nu}^{\star}\in L^{2}(\Omega\times[0,\infty);\mathsf{P}\otimes e^{-(\rho-2\lambda)t}\mathrm{d}t)$ and that $t\mapsto\mathsf{E}[|\nu^{\star}_{t}|^{2}]$ is nondecreasing, it follows that the integrability condition required in (2.19) is also met (recall that, in this section, $\vartheta_{t}\equiv\widehat{n}$). Set then $\widehat{I}^{j}:=(\hat{n},\widehat{\nu}^{j})\in\mathcal{I}_{0}$, $j\in\mathbb{N}$, and $I^{\star}:=(\hat{n},\nu^{\star})\in\mathcal{I}_{0}$.
@@
The process $\xi\in\mathcal{I}_{0}$ and it has direction of action $\hat{n}$ and intensity of action $\nu^{\xi}$ such that $\nu^{\xi}_{0^{-}}=0$ and $\mathrm{d}\nu^{\xi}_{t}=\mathrm{d}\nu^{\star}_{t}+\varepsilon e^{\lambda t}\delta(t-\tau)$, for any $t\geq 0$. Furthermore, letting $X^{x-\varepsilon\hat{n},\xi}$ be the solution to (2.7) started at time $0^{-}$ from level $x-\varepsilon\hat{n}$ and controlled by $\xi$, we have that: $X^{x-\varepsilon\hat{n},\xi}\equiv X^{x-\varepsilon\hat{n},I^{\star}}$ on $0,\tau)$, while $X^{x-\varepsilon\hat{n},\xi}\equiv X^{x,I^{\star}}$ on $\tau,\infty)$.
@@
and notice that denoting by $X^{x+\varepsilon\hat{n},\eta}$ the solution to (2.7) started at time $0^{-}$ from level $x+\varepsilon\hat{n}$ and controlled by $\eta$, we have that: $X^{x+\varepsilon\hat{n},\eta}\equiv X^{x+\varepsilon\hat{n},0}$ on $0,\tau^{\varepsilon})$, while $X^{x+\varepsilon\hat{n},\eta}\equiv X^{x,I^{\star}}$ on $\tau^{\varepsilon},\infty)$.
@@
By continuity of $U$, the stopping region $\{x\in H:\,U(x)=\Phi(x)\}$ is closed. Hence, by standard theory of optimal stopping (see, e.g., [57, Ch. I, Sec. 2.2], whose results hold for an underlying process taking values in a measurable space), one has $\mathsf{P}$-a.s. that
@@
(4.18) $U(x)=\sup_{\tau\in\mathcal{T}}\mathsf{E}\left[e^{-(\rho-\lambda)\tau}\mathds{1}_{{\tau<\theta}}\Phi(X^{x,0}_{\tau})+e^{-(\rho-\lambda)\theta}\mathds{1}_{{\tau\geq\theta}}U(X^{x,0}_{\theta})\right],\ \ \ \ \forall x\in H.$
The proof of this result does not depend on the dimension of the underlying state space and proceeds along the same lines as in the finite-dimensional case.(see, e.g., [60, Thm. 4.3] and references therein), upon exploiting the flow property of $X^{x,0}$ (see (A-6) when $I\equiv 0$).
@@
with $\mathcal{G}$ as in (3.4), and for which, recalling the class (3.2.1), we now provide the definition of viscosity solution.
@@
    We say that $u\in C(H)$ is a viscosity supersolution to (4.19) at $x\in H$ if, for every $\varphi\in\mathcal{X}$ such that $0=u(x)-\varphi(x)=\min(u-\varphi)$, one has min{((ρ-λ)-G)φ(x), (φ-Φ)(x)}0.
@@
    We say that $u\in C(H)$ is a viscosity subsolution to (4.19) at $x\in H$ if, for every $\varphi\in\mathcal{X}$ such that $0=u(x)-\varphi(x)=\max(u-\varphi)$, one has min{((ρ-λ)-G)φ(x), (φ-Φ)(x)}0.
@@
Remark 4.9.
In this case, one could also include radial functions in the definition 4.8 of viscosity solution of the considered variational inequality. We refrain from doing so for the sake of simplicity and because the goal of this section is to exploit the viscosity property – specifically, the viscosity supersolution property – using only test functions in $\mathcal{X}$ to establish a smoothness result, rather than to prepare a uniqueness result (see Remark 3.7 and compare with Remark 4.18).
@@
with $\mathcal{B}_{|\cdot|_{H}}(x,\varepsilon):=\{y\in H:\,|y-x|_{H}<\varepsilon\}.$ There exists $m_{o}>0$ such that
@@
We follow the ideas of the proof in the finite-dimensional setting proposed by [58, Thm. 5.2.1], but we simplify substantially the proof of the subsolution property thanks to Lemma 4.10.
Supersolution property. Recall (3.2.1) and let $\varphi\in\mathcal{X}$ be such that
@@
(with $\inf\emptyset=+\infty$) and letting $h>0$, by Proposition 4.7 with $\theta=\tau_{R}\wedge h$, we have
@@
By continuity, there exists some $\varepsilon>0$ such that
$((\rho-\lambda)-\mathcal{G})\varphi(y)>\eta\ \ \mbox{and}\ \ U(y)-\Phi(y)>\frac{\eta}{\rho-\lambda}\ \ \ {\forall y\in}\mathcal{B}_{|\cdot|_{H}}(x,\varepsilon),$
where $\mathcal{B}_{|\cdot|_{H}}(x,\varepsilon):=\{y\in H:\,|y-x|_{H}<\varepsilon\}.$
@@
Then, using Dynkin’s formula of Proposition 3.2 (applied again in the case of II being the null control) we obtain for each $\tau\in\mathcal{T}$
@@
If one aims at a comparison theorem for (4.19), this might be proved by adapting the techniques of [37] in order to deal with the unbounded operator in the dynamics of the state process (not present in [37]). To that end, one should treat the unbounded term as in the regular control case by adding radial functions as test functions in the definition of viscosity solutions (see [28, Ch. 3]).
@@
so that $^{G}_{\hat{n}}$ is clearly concave, belongs to $C^{1}(H)$ and it has sublinear growth with respect to $|\cdot|_{H}$. Moreover, in the first case, setting $k:=\mathcal{A}h$, one has
@@
Using (4.1) and that $|e^{t\mathcal{A}}|_{\mathcal{L}(H)}\leq e^{-\delta t}$, one easily finds from (4.15) that the semiconcavity with respect to the $|\cdot|_{-1}$ norm of $^{G}_{\hat{n}}$ implies semiconvexity of $\Phi$ with respect to the $|\cdot|_{-1}$ norm as well. Hence, given that $x\mapsto X^{x,0}_{\cdot}$ is linear, we find that $U$ is semiconvex with respect to the $|\cdot|_{-1}$ norm being supremum of semiconvex functions (see [14, Prop. 2.1.5], whose proof does not suffer the dimensionality of the state space).
By (4.15), one also has that the growth condition on $^{G}_{\hat{n}}$, together with (4.23), imply that $|\Phi(x)|\leq K_{\Phi}(1+|x|_{-1})$, for a suitable $K_{\Phi}>0$. The latter, together with (4.24), in turn gives for $\widehat{\kappa}_{o}>0$
@@
Finally, the last claim is due to [26, Cor. 2.4, p. 12] and to the fact that, being $U$ semiconvex with respect to $|\cdot|_{-1}$ norm, one can write
@@
In the rest of the paper, we are going to study the regularity properties of the (sub)gradient of $U$, that we denote by $D^{-}U$.
@@
$D^{-}U(x)\subseteq\mathcal{D}(\mathcal{A})$ at all $x\in H$.
@@
Recall that, by semiconvexity, we can write
@@
$D^{-}U(x)=D^{-}U_{0}(x)-\widehat{\kappa}_{1}{\left(\mathcal{A}^{-1}\right)^{2}}x.$
Hence, without loss of generality, up to replace $U$ by $U_{0}$, we work in the rest of this proof under the assumption that $U$ is convex.
Let $x\in H$, $(h_{n})\subseteq\mathcal{D}(\mathcal{A})$ be such that $h_{n}\to 0$ with respect to $|\cdot|_{H}$, and set $\hat{h}_{n}:=\mathcal{A}h_{n}$, so that also $|\hat{h}_{n}|_{-1}\to 0$. For $p\in D^{-}U(x)$, by convexity of $U$ we have
@@
is $|\cdot|_{H}$-upper semicontinuous by (4.25); that is, $\limsup_{n\rightarrow\infty}T_{x}(h_{n})\leq 0$. Taking now the sequence $(-h_{n})\subseteq\mathcal{D}(\mathcal{A})$ and exploiting the linearity of $T_{x}$, it also holds that $T_{x}$ is $|\cdot|_{H}$-lower semicontinuous. Hence, it is $|\cdot|_{H}$-continuous, which means that $p\in\mathcal{D}(\mathcal{A})$. ∎
Proposition 4.17.
Let $x\in H$, $\hat{h}\in K$, and set $h:=\sigma\hat{h}$. Then, $U$ is differentiable at $x$ along the direction $h$.
Proof.
We assume, without loss of generality, that $U(x)=0$ (indeed, this can be always achieved by a translation) and we notice that we can write
(4.26) $U(y)=U_{0}(y)-\frac{\widehat{\kappa}_{1}}{2}|y-x|^{2}_{-1},\quad y\in H,$
for some $\widehat{\kappa}_{1}\geq 0$ and $U_{0}$ convex.
We now argue by contradiction and assume that $U$, hence $U_{0}$, is not differentiable along the direction $h:=\sigma\hat{h}$, with $\hat{h}\in K$. By convexity of $U_{0}$, there exist $p_{1},{p}_{2}\in D^{-}U_{0}(x)$ such that
(4.27) $\langle p_{1},h\rangle_{H}\neq\langle{p}_{2},h\rangle_{H}.$
By (4.27) and since $h=\sigma\hat{h}$, we have
(4.28) $\sigma^{*}(p_{1}-p_{2})\neq 0.$
Again using convexity of $U_{0}$, we get
(4.29) $U_{0}(y)\geq\langle p_{1},y-x\rangle_{H}\vee\langle p_{2},y-x\rangle_{H},\ \ \ \forall y\in H.$
Note that, by Proposition 4.16, we have $p_{1},{p}_{2}\in\mathcal{D}(\mathcal{A})$. For $n\in\mathbb{N}$, let $\rho_{n}\colon\mathbb{R}\rightarrow\mathbb{R}$ be a function with the following properties:
(4.30a) $\rho_{n}(0)=0$
(4.30b) $\rho_{n}(t)=\rho_{n}(-t)\qquad\qquad\forall t\in\mathbb{R}$
(4.30c) $0\leq\rho_{n}(t)\leq|t|\qquad\qquad\forall t\in\mathbb{R}$
(4.30d) $\rho_{n}\in C^{2}(\mathbb{R})$
(4.30e) $\rho_{n}^{\prime}(0)=0$
(4.30f) $\rho_{n}^{\prime\prime}(0)=n$
and define $\hat{\varphi}_{n}\colon H\rightarrow\mathbb{R}$,
$\hat{\varphi}_{n}(y)\coloneqq\rho_{n}\left(\langle p_{1},y-x\rangle_{H}\vee\langle p_{2},y-x\rangle_{H}-\frac{1}{2}\langle p_{1}+p_{2},y-x\rangle_{H}\right)+\frac{1}{2}\langle p_{1}+p_{2},y-x\rangle_{H}\qquad\forall y\in H.$
Noticing that
(4.31) $\langle p_{1},y-x\rangle_{H}\vee\langle p_{2},y-x\rangle_{H}-\frac{1}{2}\langle p_{1}+p_{2},y\rangle_{H}=\frac{1}{2}\left|\langle p_{1}-p_{2},y\rangle_{H}\right|,$
we have